# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

### v1

@@ 2.3 Radial Density Profiles
The gas / clouds within a wind likely vary with projected radial distance from the launching source due to a variety of factors, including conservation of cloud surface area (e.g., 2017MNRAS.471.4061K), fountains (e.g., 2015ApJ...814...83L), or an isothermal mass-conserving gas (e.g., 2005ApJS..160..115R; 2011ApJ...734...24P).
For simplicity, gas number density $n$ can also be assumed to follow a power law such that
@@
a reasonable approximation for stellar and galactic winds (e.g., Castor1979; 1985Natur.317...44C, respectively) and a common choice in the literature (e.g., FaucherGiguere2012; 2015ApJ...801...43S; 2016MNRAS.463..541C; Flury2023; Luminari2024A&A) arising from power law solutions to differential equations (as also the case for the velocity profile) and can be indicative of momentum or energy conservation ($\beta=(\alpha-2)/2$ and $\beta=(\alpha-2)/3$, respectively FaucherGiguere2012, for the power law velocity field $\beta$). As an alternative, we can adopt an exponential profile where
@@
with a break-point or “knee” at radius $x_{1}$, inner slope $a_{1}$ and outer slope $a_{2}$. Examples of a broken power law include the shock-driven outflow launched by the interaction of winds from massive stars in proximity to Sag A ($a_{1}<a_{2}$, drop off in density profile, e.g., Quataert2004) and the relativistic outflow triggered by the waking/activating supermassive black hole in Swift J164449.3+57345 ($a_{1}>a_{2}$, flattening in density profile, e.g., Berger2012).
@@
a fast-rise exponential decay (or FRED, Equation 23, see Norris2005) distribution
@@
We show the effects of changing the density distribution for a fixed density ($n_{0}$), total column density ($N$), and velocity field ($\beta$, $v_{\infty}$) in Figure 4 assuming a spherical geometry. The emission line profile wings for the shell or pulse-like density distributions are distinct from the continuous distributions; however, distinctions among the various subsets are less clear, likely owing to the geometric dilution of the density profile by integrating over the entire volume of the line-emitting gas. The effect of pulse profiles on the emission line profile is largely consistent with early fixed-velocity single expanding shell models for Wolf-Rayet features and stellar wind bubbles wherein no line core emerges (e.g., Beals1931; TenorioTagle1996). Such marked differences in radial gas distribution may explain the line profiles of ionized gas around pre- and post-SNe stellar populations as in the different clusters in the Sunburst Arc (e.g., Mainali2022).
@@
Figure 3: Radial density profiles included in OutLines and given by Equations 18-23. Pulse-like density profiles, including the normal (‘Normal’, green, Eqn 21), log normal (‘LogNormal’, yellow, Eqn 22), and fast-rise exponential decay (‘FRED’, red, Eqn 23) distributions, represent physical scenarios such as blast waves, expanding shells, and single burst events. For visualization, the pulse-like profiles are arbitrarily set to maximize density at $r=3R_{0}$. Continuous-like density profiles, including power law (‘PowerLaw’, purple, Eqn 18), exponential (‘Exponential’, orange, Eqn 19), and double power law (‘DoublePowerLaw’, blue, Eqn 20) distributions, represent an integral over many pulses and will always maximize density at the base of the outflow where $r=R_{0}$.
Figure 4: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for fixed density, column density, and velocity field for different density distributions assuming a spherical geometry.
@@
which is likely a more physical representation of outflows in formation. We show the effects of a series of pulse-like events or damped pulses on the emergent line profiles in Figure 5 for shells resolved and unresolved in velocity. One might invoke the central limit theorem to argue that a system of many shells or bubbles will result in a Gaussian line profile; however, we find that neither the damped nor undamped pulses behave in this manner. Indeed, TenorioTagle1996 find that only a Gaussian velocity field will produce an intrinsically Gaussian line profile for a sum of many shells, a scenario which does not appear to be consistent with the bulk motion of gas predicted by theory. They further demonstrate that profiles might only appear Gaussian due to smearing by a line spread function. These results are consistent with comparisons of integrated and spatially resolved observations of the ensemble of expanding shells in 30 Doradus where the bulk velocity field produces a line profile with a Gaussian-like core but distinctly non-Gaussian wings (Chu1994).
Figure 5: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for a series of expanding bubbles, shells, or other pulse-like episodes in cases of unresolved (orange) or resolved (blue) shells with (dashed) and without (solid) damping to mimic the effects of dissipation. In all cases, density, column density, velocity field, and terminal velocity are fixed, and a spherical geometry is assumed.
@@ 2.4 Geometry
Following Flury2023, line emission or absorption is distributed evenly across the polar coordinate for each deprojected velocity (see also, e.g., Beals1931; Zheng1990; TenorioTagle1996; 2015ApJ...801...43S; Xrism2025_PDS456) allowing us to rewrite the differential solid angle as
${\rm d}\Omega={2\pi}{\rm d}w$ (27)
@@
To illustrate the effects of assumed velocity field and related cosine projection onto the line of sight, we illustrate contours of constant observed velocity (isocontours) in Figure 6 (see also Castor1970, their Figure 1, 1987ApJ...314..726L, their Figure 2). Sometimes referred to as velocity surfaces (as in Castor1970), these isocontours represent the velocities which are cosine projected to a single observed velocity, assuming a particular velocity field and varying the power law index.
@@ 2.4.1 Velocity Limits
The equal contributions of each deprojected velocity to the absorption or emission allows us to reduce the geometry to the velocity limits $w_{\ell}$ and $w_{u}$ of the integrals in Equations 5 and 10. Boundary conditions for velocity in the case of a uniform spherical outflow (illustrated in Figure 7) are given by some minimum and maximum projected velocities $w_{\ell},\ w_{u}$, respectively. The lower limit $w_{\ell}$ is simply the observed velocity $|u|$ for a uniform sphere if no obstruction occurs. From Figure 7, the line-emitting gas posterior to the source (redshifted in the case of an outflow) cannot be seen by the observer, meaning that $w_{\ell}$ must account for obstruction by the launching source at $x=1$. Solutions for the source obstruction velocity limit are given in Flury2023, which we summarize here. From the Pythagorean theorem, the minimum velocity $w_{\ell}$ must satisfy
@@
The limit of $w_{a}$ most dramatically affects the projected velocities close to the transverse axis. As a result, aperture clipping will predominantly truncate the core of the profile close to the central wavelength (see also 2015ApJ...801...43S); i.e., the broadest extent of the wings remains relatively unaffected. Moreover, absorption lines arise from gas anterior to the launch site that is backlit by the source, resulting in a “pencil beam” effect where the maximum velocity contributing to the profile is $w_{0}$ (hatches in Figure 7). Combining $w_{1}$ and $w_{0}$ gives a final upper velocity limit of
@@
Figure 8: Effect of aperture clipping on the observed line profile for nebular emission (left panel) and for a P Cygni line consisting of absorption and resonant emission with unity filling factor and $p_{res}=0.5$ (right panel). For visualization, lines are normalized to the case of no aperture clipping ($w_{a}=1$). Note that the P Cygni line simplifies to pure absorption in the case where $w_{a}=0$ ($w_{1}=w_{0}$) as in 2015ApJ...801...43S and that the effect of radiative trapping on the line core decreases as $w_{a}$ decreases.
@@ 2.4.2 Directed Outflows
Many outflows are not spherically uniform but rather are observed as bi-directional cones on a wide variety of scales (e.g., Fukui1986; Clegg1987; Heckman1990; Colina1991; Hirano1992; Marlow1995; Cameron2021). We consider a bi-directional conical geometry with a half opening angle $\theta_{o}$ and an inclination angle $i$ with respect to the observer, which we depict in Figure 10 (see also, e.g., Figure 5 in Hjelm1996). OutLines includes a tool PlotGeometry for producing such figures as a visual aid.
@@
The major difference with conical geometry is that the polar coordinate of the differential solid angle ${\rm d}\Omega$ now depends on $i$ and $\theta$; we cannot simply assume each deprojected velocity contributes equally to the integral due partial coverage. Therefore, we must account for azimuthal angle of the deprojected velocity which falls within the region of the cone projected onto the plane of the sky. The azimuthal angle inscribed by the projected cone simply scales the total solid angle by a factor $\ell_{\rm w}$$(u,w,i,\theta)\in[0,1]$ to give the relative contribution of a de-projected velocity $w$ to the observed velocity $u$. OutLines computes the conical projection from $i,\theta_{o}$ and then calculates $\ell_{\rm w}$ by determining the angular length(s) of each deprojected velocity inscribed by the projected cone. For details on this method, we refer the reader to Appendix A (see, e.g., Zheng1990; Carr2018; Luminari2018, for earlier approaches to solving this problem). We illustrate the effects of changing the opening angle $\theta_{o}$ and inclination angle $i$ in Figure 11.
Figure 11: Effects of bi-directional conical geometry on the emergent line profiles of emission (top) and absorption (bottom) profiles for fixed density (power law with $\alpha=2$) and velocity ($\beta=1$) fields. To illustrate the effects of geometry, we either vary the inclination with respect to the observer (left) from face-on ($i=0^{\circ}$) to edge-on ($i=90^{\circ}$) with fixed opening angle of $\theta=30^{\circ}$ or vary opening angle of the cones (right) from narrow ($\theta_{o}=30^{\circ}$) to a full sphere ($\theta=90^{\circ}$) with face-on inclination. The case of $i=0$ and $\theta=90^{\circ}$ (red line, right) is consistent with a filled sphere. For visualization, line profiles are all normalized to the spherical case ($i=0^{\circ}$, $\theta=90^{\circ}$).
@@ 2.4.3 Directed Outflows with Cavities
An additional geometry observed in young stellar objects (e.g., Padgett1999; Louvet2018; deValon2020), stellar winds (e.g. Sirianni1998), starbursts (e.g., Axon1978; Marlow1995; Heckman1995; Rich2010; Yoshida2010), and AGN (e.g., Phillips1983; Hjelm1996; Crenshaw2000a; Crenshaw2000b; Muller2011; Venturi2018) is that of hollow cones of outflowing material. We include this geometry, assuming some cavity angle $\theta_{c}$ and opening angle $\theta_{o}$ (see also models by Das2006; Muller2011; Bae2016; Luminari2024A&A). OutLines implements this geometry by subtracting the fractional angular length $\ell_{\rm cavity}$ of each deprojected velocity inscribed by the cavity from fractional the angular length inscribed by the total cone (see Appendix A). The difference will always be positive and greater than zero since $\theta_{o}>\theta_{c}$ by definition. We illustrate the impact of a hollowing of the outflow cone on the emergent line profile in Figure 12. Somewhat intuitively, an edge-on hollow-cone outflow loses flux from the least Doppler-shifted gas (truncating the core) while a face-on hollow-cone outflow loses flux from the most Doppler-shifted gas (truncating the wings).
@@ 2.4.4 An Obstructing Disk
The outflow may be launched from a region embedded in a disk of optically thick material, a phenomenon observed in a variety of astrophysical scenarios: expelled debris in planetary nebulae (e.g., Icke1981), protoplanetary material around young stellar objects (e.g., Padgett1999; Ilee2016), the dusty material concentrated around an AGN (e.g., Veilleux2001), or even the plane of a galaxy (e.g., Hjelm1996; Cronin2025). In such cases, the disk will obstruct some of the observer’s view of the outflow. OutLines optionally includes obstruction by a circular disk which is orthogonal to the central axis of the cones and which has some radius $x_{d}=r_{\rm disk}/R_{0}$. As with the cones, the disk must be projected onto the plane of the sky and subsequently compared to the deprojected velocity inscribed by the projected cone to determine the fractional arc length(s) $\ell_{\rm w}$ inscribed by the cone but not the projected disk. We elaborate on this formalism in Appendix A. Because a dense disk of material will likely suppress any outwardly moving material within its plane, OutLines only implements a disk if cones are present.
@@
We illustrate the effects of including an obstructing disk in Figure 13. By definition, the disk will only obstruct the posterior cone, so if $i+\theta_{o}<90^{\circ}$, the disk will only obstruct the receding outflow, solely obstructing the redshifted component (left of Figure 13). However, when $i+\theta_{o}>90^{\circ}$, the disk obstructs the lune projected by the posterior cone toward the observer, resulting in loss of the blueshifted component of the line profile as well (right of Figure 13).

### v2

@@ 2.3 Radial Density Profiles
The gas / clouds within a wind likely vary with projected radial distance from the launching source due to a variety of factors, including conservation of cloud surface area (e.g., Krumholz et al., 2017), fountains (e.g., Leroy et al., 2015), or an isothermal mass-conserving gas (e.g., Rupke et al., 2005b; Prochaska et al., 2011).
We adopt a smooth function to describe a continuous distribution of delta functions, with each delta function “pulse” representing an instantaneous burst or ejection of material with gas density $n$ from the launch site and projected radially. For simplicity, $n$ can be assumed to follow a power law such that
@@
a reasonable approximation for stellar and galactic winds (e.g., Castor and Lamers, 1979; Chevalier and Clegg, 1985, respectively) and a common choice in the literature (e.g., Faucher-Giguère and Quataert, 2012; Scarlata and Panagia, 2015; Chisholm et al., 2016; Flury et al., 2023; Luminari et al., 2024) arising from power law solutions to differential equations (as also the case for the velocity profile) and can be indicative of momentum or energy conservation ($\beta=(\alpha-2)/2$ and $\beta=(\alpha-2)/3$, respectively Faucher-Giguère and Quataert, 2012, for the power law velocity field $\beta$). Perhaps intuitively, $\alpha=2$ represents a mass-conserving scenario with density changing only as a $r^{-2}$ projection effect. As an alternative, we can adopt an exponential profile where
@@
with a break-point or “knee” at radius $x_{1}$, inner slope $a_{1}$ and outer slope $a_{2}$. Examples of a broken power law include the shock-driven outflow launched by the interaction of winds from massive stars in proximity to Sag A ($a_{1}<a_{2}$, drop off in density profile, e.g., Quataert, 2004) and the relativistic outflow triggered by the waking/activating supermassive black hole in Swift J164449.3+57345 ($a_{1}>a_{2}$, flattening in density profile, e.g., Berger et al., 2012).
@@
a fast-rise exponential decay (or FRED, Equation 24, see Norris et al. 2005) distribution
@@
a logistic column density distribution
$n[x]=\frac{k\exp[-k(x-x0)]}{(1+\exp[-k(x-x0)])^{2}}{\rm,\hskip 10.00002ptDLogistic,}$ (25)
@@
We show the effects of changing the density distribution for a fixed density ($n_{0}$), total column density ($N$), and velocity field ($\beta$, $v_{\infty}$) in Figure 4 assuming a spherical geometry. The emission line profile wings for the shell or pulse-like density distributions are distinct from the continuous distributions; however, distinctions among the various subsets are less clear, likely owing to the geometric dilution of the density profile by integrating over the entire volume of the line-emitting gas. The effect of pulse profiles on the emission line profile is largely consistent with early fixed-velocity single expanding shell models for Wolf-Rayet features and stellar wind bubbles wherein no line core emerges (e.g., Beals, 1931; Tenorio-Tagle et al., 1996). Such marked differences in radial gas distribution may explain the line profiles of ionized gas around pre- and post-SNe stellar populations as in the different clusters in the Sunburst Arc (e.g., Mainali et al., 2022).
@@
Figure 3: Radial density profiles included in OutLines and given by Equations 19-28. Top: Wind-like density profiles, including power law (‘PowerLaw’, purple, Eqn 19), double power law (‘DoublePowerLaw’, yellow, Eqn 21), and exponential (‘Exponential’, green, Eqn 20) distributions. Center: Pulse-like density profiles, including the normal (‘Normal’, blue, Eqn 22), log normal (‘LogNormal’, red, Eqn 23), fast-rise exponential decay (‘FRED’, orange, Eqn 24), and derivative of the logistic (‘DLogic’, indigo, Eqn 25) distributions, represent physical scenarios such as blast waves, expanding shells, and single burst events. Bottom: Ensembles of pulse-like profiles without (black solid, Eqn 27) and with (black dashed, Eqn 28) damping. These ensembles represent the resolved elements of a radially clumpy wind or a series of outflowing episodes.
Figure 4: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for fixed flux, column density, and velocity field for different density distributions assuming a spherical geometry. Colors matched to the corresponding profiles in Figure 3.
@@
which is likely a more physical representation of outflows in formation. We show the effects of a series of pulse-like events or damped pulses on the emergent line profiles in Figure 5 for shells resolved and unresolved in velocity. Figure 5 illustrates how infinitely thin shells with infinitely small separations ($\sigma_{x},x_{k}\to 0$) or relatively thick shells ($\sigma_{k}\gtrsim x_{k}$) produce a wind-like profile.
One might invoke the central limit theorem to argue that a system of many shells or bubbles will result in a Gaussian line profile; however, we find that neither the damped nor undamped pulses behave in this manner. Indeed, Tenorio-Tagle et al. (1996) find that only a Gaussian velocity field will produce an intrinsically Gaussian line profile for a sum of many shells, a scenario which does not appear to be consistent with the bulk motion of gas predicted by theory. They further demonstrate that profiles might only appear Gaussian due to smearing by a line spread function. These results are consistent with comparisons of integrated and spatially resolved observations of the ensemble of expanding shells in 30 Doradus where the bulk velocity field produces a line profile with a Gaussian-like core but distinctly non-Gaussian wings (Chu and Kennicutt, 1994).
Figure 5: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for an ensemble of expanding bubbles, shells, or other pulse-like episodes in cases of thick (orange, $\sigma_{x}=0.5$) or thin (blue, $\sigma_{x}=0.1$) widely-spaced shells and thin closely-spaced shells (green, $x_{k}=0.2$) with (dashed) and without (solid) damping to mimic the effects of dissipation. In all cases, density, column density, velocity field, and terminal velocity are fixed, and a spherical geometry is assumed.
@@ 2.4 Geometry
For any given radial velocity $w$, line emission or absorption is distributed evenly across a given projection angle $\varrho$ (the polar coordinate), which we illustrate here under full special relativistic treatment. The special relativistic velocity projection of $w$ in an arbitrary polar direction $\varrho^{\prime}$ in the observer’s frame incorporates both longitudinal and transverse effects such that, following Einstein (1905, their §7),
$\lambda=\lambda_{0}\gamma\left(1+\omega\cos\varrho^{\prime}\right)$ (29)
where $\omega=w\>v_{\infty}/c$ and $\gamma^{-1}=\sqrt{1-\omega^{2}}$ is the Lorenzt factor. Additionally from §7 of Einstein (1905), the angle $\varrho$ in the source frame requires a relativistic aberration correction to the observer’s frame $\varrho^{\prime}$, with a differential which reduces to
$\sin\varrho^{\prime}{\rm d}\varrho^{\prime}=\left(\frac{\lambda_{0}}{\lambda}\right)^{2}\sin\varrho{\rm d}\varrho.$ (30)
Combining these identities and respective differentials with Equation 1 provides a convenient substitution for the $\sin\varrho{\rm d}\varrho$ term in ${\rm d}\Omega$ such that
${\rm d}\Omega^{\prime}=2\pi\frac{\sqrt{1-\omega^{2}}}{\omega(1-\omega)^{2}}\sqrt{\frac{1-\omega}{1+\omega}}{\rm d}\omega$ (31)
@@
To illustrate the effects of assumed velocity field and related $\cos\varrho$ projection onto the line of sight, we illustrate contours of constant observed velocity (isocontours) in Figure 6 (see also Castor 1970, their Figure 1, Lucy, 1971, their Figure 2). Sometimes referred to as velocity surfaces (as in Castor, 1970), these isocontours represent the velocities which are cosine projected to a single observed velocity, assuming a particular velocity field and varying the power law index.
@@ 2.4.1 Velocity Limits
The weighted contributions of each deprojected velocity to the absorption or emission allows us to reduce the geometry to the velocity limits $w_{\ell}$ and $w_{u}$ of the integrals in Equations 5 and 10. Boundary conditions for velocity in the case of a uniform spherical outflow (illustrated in Figure 7) are given by some minimum and maximum projected velocities $w_{\ell},\ w_{u}$, respectively. The lower limit $w_{\ell}$ is simply the observed velocity $|u|$ for a uniform sphere if no obstruction occurs. From Figure 7, the line-emitting gas posterior to the source (redshifted in the case of an outflow) cannot be seen by the observer, meaning that $w_{\ell}$ must account for obstruction by the launching source at $x=1$. Solutions for the source obstruction velocity limit are given in Flury et al. (2023), which we summarize here. From the Pythagorean theorem, the minimum velocity $w_{\ell}$ must satisfy
@@
The limit of $w_{a}$ most dramatically affects the projected velocities close to the transverse axis. As a result, aperture clipping will predominantly truncate the core of the profile close to the central wavelength (see also Scarlata and Panagia, 2015); i.e., the broadest extent of the wings remains relatively unaffected. Moreover, absorption lines arise from gas anterior to the launch site that is backlit by the source, resulting in a “pencil beam” effect where the maximum velocity contributing to the profile is $w_{0}$ (hatches in Figure 7). Combining $w_{1}$ and $w_{0}$ gives a final upper velocity limit of
@@
Figure 8: Effect of aperture clipping on the observed line profile for a P Cygni line consisting of absorption and resonant emission with filling factor $\epsilon=0.2$ and escape fraction $p_{res}=1$ (left panel) and for nebular emission (right panel). For visualization, lines are normalized to the case of no aperture clipping ($w_{a}=1$). Dotted black line in the left panel indicates pure absorption, demonstrating that infilling can still occur in the fully aperture-clipped limit but that only an absorption profile is observed (cf. Scarlata and Panagia, 2015). Note that the effect of radiative trapping on the resonant line core decreases as $w_{a}$ decreases.
@@ 2.4.2 Directed Outflows
Many outflows are not spherically uniform but rather are observed as bi-directional cones on a wide variety of scales (e.g., Fukui et al., 1986; Clegg et al., 1987; Heckman et al., 1990; Colina et al., 1991; Hirano et al., 1992; Marlowe et al., 1995; Cameron et al., 2021). We consider a bi-directional conical geometry with a half opening angle $\theta_{o}$ and an inclination angle $i$ with respect to the observer, which we depict in Figure 10 (see also, e.g., Figure 5 in Hjelm and Lindblad 1996). OutLines includes a tool PlotGeometry for producing such figures as a visual aid.
@@
The major difference with conical geometry is that the polar coordinate of the differential solid angle ${\rm d}\Omega$ now depends on $i$ and $\theta$; we cannot simply assume each deprojected velocity contributes equally to the integral due partial coverage. Therefore, we must account for azimuthal angle of the deprojected velocity which falls within the region of the cone projected onto the plane of the sky. The azimuthal angle inscribed by the projected cone simply scales the total solid angle by a factor $\ell_{\rm w}$$(u,w,i,\theta)\in[0,1]$ to give the relative contribution of a de-projected velocity $w$ to the observed velocity $u$. OutLines computes the conical projection from $i,\theta_{o}$ and then calculates $\ell_{\rm w}$ by determining the angular length(s) of each deprojected velocity inscribed by the projected cone. For details on this method, we refer the reader to Appendix A (see, e.g., Zheng et al., 1990; Carr et al., 2018; Luminari et al., 2018, for earlier approaches to solving this problem). We illustrate the effects of changing the opening angle $\theta_{o}$ and inclination angle $i$ in Figure 11.
Figure 11: Effects of bi-directional conical geometry on the emergent line profiles of emission (top) and absorption (bottom) profiles for fixed density (power law with $\alpha=2$) and velocity ($\beta=1$) fields. To illustrate the effects of geometry, we either vary the inclination with respect to the observer (left) from face-on ($i=0^{\circ}$) to edge-on ($i=90^{\circ}$) with fixed opening angle of $\theta=30^{\circ}$ or vary opening angle of the cones (right) from narrow ($\theta_{o}=30^{\circ}$) to a full sphere ($\theta=90^{\circ}$) with face-on inclination. The case of $i=0$ and $\theta=90^{\circ}$ (red line, right) is consistent with a filled sphere. For visualization, line profiles are all normalized to the spherical case ($i=0^{\circ}$, $\theta=90^{\circ}$). The effects of inclination on the emergent line profile shown in the upper left are similar to the 3D Monte Carlo simulation predictions for X-ray fluorescent lines from the skirt code (Vander Meulen et al., 2024; Bianchi et al., 2026), with face-on inclination yielding broader and flatter profiles and edge-on inclination yielding narrower and peakier profiles.
@@ 2.4.3 Directed Outflows with Cavities
An additional geometry observed in young stellar objects (e.g., Padgett et al., 1999; Louvet et al., 2018; de Valon et al., 2020), stellar winds (e.g. Sirianni et al., 1998), starbursts (e.g., Axon and Taylor, 1978; Marlowe et al., 1995; Heckman et al., 1995; Rich et al., 2010; Yoshida et al., 2011), and AGN (e.g., Phillips et al., 1983; Hjelm and Lindblad, 1996; Crenshaw and Kraemer, 2000; Crenshaw et al., 2000; Müller-Sánchez et al., 2011; Venturi et al., 2018) is that of hollow cones of outflowing material. We include this geometry, assuming some cavity angle $\theta_{c}$ and opening angle $\theta_{o}$ (see also models by Das et al., 2006; Müller-Sánchez et al., 2011; Bae and Woo, 2016; Luminari et al., 2024). OutLines implements this geometry by subtracting the fractional angular length $\ell_{\rm cavity}$ of each deprojected velocity inscribed by the cavity from fractional the angular length inscribed by the total cone (see Appendix A). The difference will always be positive and greater than zero since $\theta_{o}>\theta_{c}$ by definition. We illustrate the impact of a hollowing of the outflow cone on the emergent line profile in Figure 12. Somewhat intuitively, an edge-on hollow-cone outflow loses flux from the least Doppler-shifted gas (truncating the core) while a face-on hollow-cone outflow loses flux from the most Doppler-shifted gas (truncating the wings).
@@ 2.4.4 An Obstructing Disk
The outflow may be launched from a region embedded in a disk of optically thick material, a phenomenon observed in a variety of astrophysical scenarios: expelled debris in planetary nebulae (e.g., Icke, 1981), protoplanetary material around young stellar objects (e.g., Padgett et al., 1999; Ilee et al., 2016), the dusty material concentrated around an AGN (e.g., Veilleux et al., 2001), or even the plane of a galaxy (e.g., Hjelm and Lindblad, 1996; Cronin et al., 2025). In such cases, the disk will obstruct some of the observer’s view of the outflow. OutLines optionally includes obstruction by a circular disk which is orthogonal to the central axis of the cones and which has some radius $x_{d}=r_{\rm disk}/R_{0}$. As with the cones, the disk must be projected onto the plane of the sky and subsequently compared to the deprojected velocity inscribed by the projected cone to determine the fractional arc length(s) $\ell_{\rm w}$ inscribed by the cone but not the projected disk. We elaborate on this formalism in Appendix A. OutLines implements a disk either in tandem with the cone geometries or alone as a hemisphere geometry.
Figure 13: Same as Figure 11 but for the effects of an obstructing disk as a function of inclination with respect to the observer in (left) a thin single-shell hemisphere geometry where all material posterior to the disk is obstructed (cf. Xrism Collaboration et al., 2025, their Fig 12) and (right) a biconical wind with $\theta_{0}=30^{\circ}$ and $x_{\rm disk}=5$. Colors indicate changes to the inclination angle $i$, illustrating the redistribution of flux from entirely blue-shifted ($i=0^{\circ}$) to evenly blue- and red-shifted ($i=90^{\circ}$).
@@
We illustrate the effects of including an obstructing disk for the hemisphere and bidirectional cone cases in Figure 13 and for the bidirectional cones in Figure 14. By definition, the disk will only obstruct the posterior cone, so if $i+\theta_{o}<90^{\circ}$, the disk will only obstruct the receding outflow, solely obstructing the redshifted component (left of Figure 14). However, when $i+\theta_{o}>90^{\circ}$, the disk obstructs the lune projected by the posterior cone toward the observer, resulting in loss of the blueshifted component of the line profile as well (right of Figure 14).